OpenStudy (anonymous):

What is the second step to solve the equation 2b + 5 = negative 17

OpenStudy (anonymous):

2b+5=-17 2b +5 -5= -17 -5 2b=-22 2b / 2 = -22 / 2 b= -11

OpenStudy (anonymous):

Thanks!